OpenStudy (anonymous):

The distance traveled at a constant speed is directly proportional to the time of travel. If Olivia traveled 306 miles in 8.5 hours, how many miles will Olivia travel in 9.1 hours at the same constant speed?

hartnn (hartnn):

let Olivia travel x miles. distance1/time1=distance2/time2 so, 306/8.5=x/9.1 so x= (306/8.5)*9.1 =??

OpenStudy (anonymous):

327.6 miles